OpenStudy (anonymous):

Is y=4x+5 parallel or perpendicular or neither

OpenStudy (anonymous):

it is parllel

OpenStudy (anonymous):

when u talk about a line being parallel or perpendicular or intersecting then it is with respect to any other line or any axis.

OpenStudy (anonymous):

It is neither parallel nor is it perpendicular because those are relationships between two lines, and y=4x+5 is only itself a single line.

OpenStudy (doc.brown):

Parallel or perpendicular to what? It's a line.

OpenStudy (unklerhaukus):

y=4x+5 is parallel to y=4x+b and perpendicular to y=-x/4+b where b is some y-intercept